OpenStudy (anonymous):

use a table to find the solution of the system y=7-x y=x+1

OpenStudy (anonymous):

can you rewrite the second equation so that x is isolated?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

ok, let me help. so since you have y=7-x, you want to carry the x over to the left and the y over to the right. add x to both sides and subtract y from both sides getting: x=-y+7

OpenStudy (anonymous):

thats the answer?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

but you plug it into the second equation

OpenStudy (anonymous):

try plugging x=-y+7 into y=x+1

OpenStudy (anonymous):

what is it

OpenStudy (anonymous):

can you try it?

OpenStudy (anonymous):